OpenStudy (anonymous):

i'm stuck i need help factoring out this> 7k^2+9k

OpenStudy (anonymous):

each has a common factor of \(k\) so you can "factor it out"

OpenStudy (anonymous):

so i keep getting 16 but i'm not sure if thats right

OpenStudy (callisto):

It's not correct. When you take out the common factor k, what is left for each term?

OpenStudy (anonymous):

7 and 9

OpenStudy (lgbasallote):

remember.. \[7k^2 = 7k \times k\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i have 63

OpenStudy (callisto):

You cannot multiply the terms... One example: factor x^2 + 2x In this case, x is the common factor. Take it out and group the rest of the terms, that is x^2 + 2x = x (x) + 2(x) = x ( x+2) Can you try again for your question?

OpenStudy (anonymous):

so is k(7k+9) correct

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

thanks